### v1

# Universal $L_{2}$-approximation using median lattice algorithms

Abstract
We study the problem of multivariate $L_{2}$-approximation of functions in a weighted Korobov space using a median lattice-based algorithm recently proposed by the authors. In the original work, the algorithm requires knowledge of the smoothness and weights of the Korobov space to construct the hyperbolic cross index set, where each coefficient is estimated via the median of approximations obtained from randomly shifted, randomly chosen rank-1 lattice rules. In this paper, we introduce a universal median lattice-based algorithm, which eliminates the need for any prior information on smoothness and weights. Although the tractability property of the algorithm slightly deteriorates, we prove that, for individual functions in the Korobov space with arbitrary smoothness and (downward-closed) weights, it achieves an $L_{2}$-approximation error arbitrarily close to the optimal rate with respect to the number of function evaluations.
Keywords: multivariate $L_{2}$-approximation, median algorithm, rank-1 lattice rule, weighted Korobov space, universality
AMS subject classifications: 41A25, 41A63, 65D15, 65D30, 65Y20
1 Introduction
This paper deals with the classical problem of approximating functions in the $L_{2}$-norm, defined on a so-called weighted Korobov space, which consists of $d$-variate, one-periodic functions with sufficiently fast decaying Fourier coefficients. Korobov spaces are reproducing kernel Hilbert spaces, and it is known that their elements can be represented pointwise by absolutely convergent Fourier series, i.e., we are dealing with functions of the form
$f(\boldsymbol{x})=\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\widehat{f}(\boldsymbol{h})\exp(2\pi\mathtt{i}\boldsymbol{h}\cdot\boldsymbol{x}),$
where, as usual,
$\widehat{f}({\boldsymbol{h}}):=\int_{[0,1)^{d}}f(\boldsymbol{x})\exp({-2\pi\mathtt{i}\boldsymbol{h}\cdot\boldsymbol{x}})\,\mathrm{d}\boldsymbol{x}$
is the $\boldsymbol{h}$-th Fourier coefficient of $f$ for $\boldsymbol{h}\in\mathbb{Z}^{d}$, where $\mathbb{Z}$ denotes the set of integers.
Korobov spaces are of particular interest in the context of quasi-Monte Carlo (QMC) rules, and have been studied frequently in the literature, see, e.g., [8, 9, 21] and the references therein. Indeed, we shall also use QMC methods in the present paper; these methods are equal-weight quadrature rules that are frequently used for the numerical integration of multivariate functions. To be more precise, QMC rules have the form
$\frac{1}{N}\sum_{k=0}^{N-1}g(\boldsymbol{x}_{k})\approx\int_{[0,1]^{d}}g(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}$
for a suitably defined $g$, where $\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N-1}$ are carefully selected sampling points.
In the context of $L_{2}$-approximation, QMC rules are typically employed for approximately computing suitably selected Fourier coefficients of the function to be approximated, or in splines. Previous work analyzing such methods include [1, 3, 5, 18, 19, 22, 27] and many others. The approach considered here uses rank-1 lattice point sets as the QMC samples, and in particular studies the use of so-called median rules. Median rules have first been dealt with for numerical integration and are based on the following principle. When working with QMC integration rules, one often introduces a suitable randomization thereof, and then repeats the integration algorithm for independent realizations of the randomized rule. For median rules, one then takes the median of these independent instances as an approximation to an integral. A considerable advantage of median rules over usual QMC rules is that they are frequently independent of the parameters (as, e.g., the smoothness) of particular function spaces, and therefore offer a high degree of universality. Median rules for numerical integration have been presented, e.g., in [11, 12, 23]. In the recent paper [22], the authors have successfully used median lattice algorithms for $L_{2}$-approximation in weighted Korobov spaces. However, a drawback of the results in that paper is the partial loss of the universality of the median approach, due to the definition of the algorithm in [22]. The goal of the present paper is to provide an alternative algorithm that regains this universality. As we shall see below, our algorithm can be shown to converge at a rate that is arbitrarily close to optimal. By “arbitrarily close to optimal” we mean that if we consider the number $M$ of function evaluations of a given function $f$ required by our algorithm to compute an approximation to $f$, the algorithm converges with a rate of $\mathcal{O}(M^{-\alpha+\varepsilon})$ for arbitrarily small $\varepsilon>0$, where $\alpha$ is the smoothness parameter of the Korobov space under consideration (see below for the precise definition of Korobov spaces).
As shown in [18] and related works, using a single lattice rule with $M$ nodes to approximate the Fourier coefficients yields a convergence rate of the worst-case (i.e., considering the supremum over all functions in the unit ball of the space) $L_{2}$-approximation error (see Section 2) arbitrarily close to $\mathcal{O}(M^{-\alpha/2})$. This result is in a certain sense best possible, because, in a slightly more general setting, [2] proves that the rate $\mathcal{O}(M^{-\alpha/2})$ cannot be improved with such an approach for any $d\geq 2$. However, as shown in [13] and [14], this convergence rate can be improved in a probabilistic sense by employing so-called multiple lattices. Moreover, in the recent paper [3] by Cai, Goda, and Kazashi, the authors introduced a randomized lattice-based algorithm for $L_{2}$-approximation and proved that one can achieve a convergence rate arbitrarily close to $\mathcal{O}(M^{-\alpha(2\alpha+1)/(4\alpha+1)+\varepsilon})$ of the worst-case root-mean-squared $L_{2}$-approximation error, which is better than the rate $\mathcal{O}(M^{-\alpha/2})$. The same paper proves that this randomized algorithm cannot achieve a better rate than $\mathcal{O}(M^{-\alpha/2-1/2})$. In [22], we used median lattice rules based on $R$ randomized lattice point sets with $N$ points each, i.e., our algorithm required $M=RN$ function values of a function $f$ to be approximated. For such algorithms, we could show that we achieve a root-mean-squared error with a convergence order arbitrarily close to $\mathcal{O}(M^{-\alpha})$, and this result is best possible for Korobov spaces. However, as pointed out above, the drawback of the result in [22] is that the algorithm is not universal with respect to the parameters of the Korobov space, which we try to overcome in the present paper. (Note that there exist other approaches by various authors that yield a convergence rate arbitrarily close to $\mathcal{O}(M^{-\alpha})$, but these are not based on lattice rules or QMC methods; see, e.g., [4, 10, 15, 26].)
Moreover, we will show that, under suitable assumptions, this convergence can take place with only a mild dependence on the dimension of the problem. We will achieve this by making use of the concept of weighted function spaces in the sense of Sloan and Woźniakowski, see [25]. To this end, let $\boldsymbol{\gamma}=(\gamma_{\mathfrak{u}})_{\mathfrak{u}\subseteq\{1{:}d\}}$ be a collection of weights. Here and in the following, we write $\{1{:}V\}$ to denote the set $\{1,\ldots,V\}$ for a positive integer $V$. The basic idea of weights is to assign a positive number to every possible group of variables $\mathfrak{u}\subseteq\{1{:}d\}$, which describes how much influence the respective groups of variables have in the problem. Large values of $\gamma_{\mathfrak{u}}$ indicate higher influence, low values of $\gamma_{\mathfrak{u}}$ indicate less influence. These weights are incorporated in the inner product and norm of the weighted Korobov space. It is known that under certain summability conditions on the weights the curse of dimensionality can be weakened or even vanquished completely. This corresponds to the idea that, while the approximation problem may nominally depend on a large number of variables, their influence on the problem diminishes fast, such that only rather few of them contribute significantly to the problem.
The rest of the paper is structured as follows. In Section 2, we introduce the necessary preliminaries and notation, including the weighted Korobov space, the $L_{2}$-approximation problem, and rank-1 lattice rules. Section 3 presents the universal median lattice-based algorithm and states the main probabilistic $L_{2}$-error bound (Theorem 3.8), along with a discussion of its tractability properties. In Section 4, we provide a detailed analysis of the three terms in the error decomposition ($\clubsuit$3.2), establishing probabilistic bounds that, when combined, yield the main result. We conclude this paper with a discussion in Section 5.
2 Preliminaries and notation
In what follows, we denote the set of integers by $\mathbb{Z}$ and the set of positive integers by $\mathbb{N}$. As usual, we denote the exponential function by $\exp$, and write $\exp(1)=e$ for short. Bold symbols are used to denote vectors, whose lengths are usually clear from the context. We denote the imaginary unit by $\mathtt{i}:=\sqrt{-1}$, and further notation will be introduced as necessary.
2.1 The weighted Korobov space
Let $\boldsymbol{\gamma}=(\gamma_{\mathfrak{u}})_{\mathfrak{u}\subseteq\{1{:}d\}}$ be a collection of positive real numbers, which we refer to as the weights in what follows. As mentioned in the introduction, the weight $\gamma_{\mathfrak{u}}$ models the influence of the group of variables $x_{j}$ with indices $j\in\mathfrak{u}$. Throughout this paper, we assume $0<\gamma_{\mathfrak{u}}\leq 1$; this assumption could be relaxed without changing the essence of our results, at the cost of additional technical notation. We further assume that the weights satisfy $\gamma_{\mathfrak{v}}\leq\gamma_{\mathfrak{w}}$ whenever $\mathfrak{w}\subseteq\mathfrak{v}$, a property we refer to as being downward-closed.
For a collection of weights $\boldsymbol{\gamma}=(\gamma_{\mathfrak{u}})_{\mathfrak{u}\subseteq\{1{:}d\}}$ and a real number $\alpha>1/2$, we define, for $\boldsymbol{h}=(h_{1},\ldots,h_{d})\in\mathbb{Z}^{d}$,
$r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}):=\gamma_{\operatorname{supp}(\boldsymbol{h})}^{-1}\prod_{j\in\operatorname{supp}(\boldsymbol{h})}\left|h_{j}\right|^{2\alpha},$
where $\operatorname{supp}(\boldsymbol{h})=\{j\in\{1{:}d\}\,\colon\,h_{j}\neq 0\}$.
We now define the weighted Korobov space $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, which is a reproducing kernel Hilbert space consisting of one-periodic functions on $0,1)^{d}$, where periodicity is understood with respect to each variable. The norm in $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$ is defined in terms of the Fourier coefficients by
$\|f\|_{d,\alpha,\boldsymbol{\gamma}}^{2}:=\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}\left|\widehat{f}({\boldsymbol{h}})\right|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}).$
Note that $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$ is a subspace of $L_{2}([0,1)^{d})$, and that all elements of the Korobov space with $\alpha>1/2$ can be represented pointwise by their Fourier series. Moreover, it is known that if $\alpha\in\mathbb{N}$, the univariate Korobov (periodic Sobolev) space that serves as the building block of $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$ consists precisely of one-periodic functions on $[0,1)$ whose derivatives up to order $\alpha-1$ are absolutely continuous and whose $\alpha$-th derivative belongs to $L_{2}([0,1))$. For this reason, $\alpha$ is called the smoothness parameter. In the classical product-weight setting (i.e., when $\gamma_{\mathfrak{u}}=\prod_{j\in\mathfrak{u}}\gamma_{j}$), the $d$-dimensional space $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$ is the $d$-fold tensor product of these univariate spaces. For general weights $\boldsymbol{\gamma}$, the space is defined by the Fourier-weighted norm above; see [7] for further details.
2.2 $L_{2}$-approximation
We will study $L_{2}$-approximation of functions $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$. More formally, this means that we study the approximation of the embedding operator $S:\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}\to L_{2}([0,1)^{d})$, $S(f)=f$. For a given deterministic approximation algorithm $A:\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}\to L_{2}([0,1)^{d})$, a common error measure considered in the literature is the worst-case error,
$\mathrm{err}(\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}},L_{2},A):=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}\\ \left\|f\right\|_{d,\alpha,\boldsymbol{\gamma}}\leq 1\end{subarray}}\|f-A(f)\|_{L_{2}},$
i.e., one considers the worst performance of $A$ across the unit ball of $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$.
In this paper, we will deal with randomized algorithms instead of deterministic algorithms. When we speak of a randomized approximation algorithm, we refer to a pair consisting of a probability space $(\Omega,\Sigma,\mu)$ and a family of mappings $A=(A^{\omega})_{\omega\in\Omega}$, where each $A^{\omega}$ is a deterministic approximation algorithm for fixed $\omega\in\Omega$. In the randomized setting, a common error measure is the worst-case root-mean-squared error, also referred to as the randomized error, defined by
$\mathrm{err}^{\mathrm{ran}}(\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}},L_{2},(A^{\omega})):=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}\\ \left\|f\right\|_{d,\alpha,\boldsymbol{\gamma}}\leq 1\end{subarray}}\left(\mathbb{E}_{\omega}\left[\|f-A^{\omega}(f)\|^{2}_{L_{2}}\right]\right)^{1/2}.$
Alternatively, one can consider the $(\epsilon,\delta)$-approximation framework, see for instance [17], which asks for which pairs $(\epsilon,\delta)$ the algorithm satisfies
$\mathrm{Pr}_{\omega}\left[\|f-A^{\omega}(f)\|_{L_{2}}>\epsilon\right]\leq\delta\hskip 18.49988pt\text{for all $f$ with $\left\|f\right\|_{d,\alpha,\boldsymbol{\gamma}}\leq 1$.}$ (1)
Our main result (see Corollary 3.8 below) is that we can reach a probabilistic error bound of order $M^{-\alpha}$ when using $M$ function evaluations with high probability, by using a median algorithm (see Section 3 for the precise definition), where the median is taken over $R$ randomized instances of QMC rules. Here, $M$ is of the form $RN$, where $N$ is the number of integration nodes used in each single QMC rule. The essential progress in comparison to our previous result in [22] is that the algorithm will not be depending on the parameters $\alpha$ and $\boldsymbol{\gamma}$ of $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, and thus can be considered as much more universal.
2.3 Rank-1 lattice rules
Let us now outline which type of QMC rules we use in our median algorithm, which are rank-1 lattice rules. For simplicity, let us assume that $N$ is a prime in the following (assuming that $N$ is prime guarantees that all one-dimensional projections of a lattice point set are evenly distributed, which is an advantage. If we would allow composite $N$, we presumably could derive similar results, but would need more technical notation). Let $\boldsymbol{z}=(z_{1},\ldots,z_{d})$ be a vector with each component in $\{1{:}(N-1)\}$. We can then define a lattice point set $\mathcal{P}_{N}$ with points $\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N-1}$ as follows. For $j\in\{1{:}d\}$ and $k\in\{0,\ldots,N-1\}$, the $j$-th component of $\boldsymbol{x}_{k}$ is given by
$x_{k}^{(j)}:=\left\{\frac{kz_{j}}{N}\right\},$
where $\{y\}=y-\lfloor y\rfloor$ denotes the fractional part of a real number $y$. By repeating the above procedure for all $j\in\{1{:}d\}$ and $k\in\{0,\ldots,N-1\}$, we obtain the full lattice point set $\mathcal{P}_{N}$. A QMC rule using $\mathcal{P}_{N}$ is called a (rank-1) lattice rule.
Note that, for fixed $N$ and $d$, a lattice rule is fully characterized by the choice of the generating vector $\boldsymbol{z}$. Not all choices of $\boldsymbol{z}$ yield lattice rules that have sufficient quality to be usefully employed in integration or approximation algorithms. However, there are fast construction algorithms available that return good generating vectors for given $N$, $d$, $\alpha$, and $\boldsymbol{\gamma}$. We refer to the books [8, 24] for overviews of the theory of lattice rules, and in particular to [8, Chapters 3 and 4] for constructions of good lattice rules. In the present paper, however, we will make a random choice of the generating vectors $\boldsymbol{z}$ to obtain the lattice point sets used. Therefore, we need not be concerned with construction algorithms here, which is a general computational advantage of median rules.
It is sometimes useful, and we shall also do so here, to introduce an additional random element when applying lattice rules. This is commonly achieved by a random shift, $\Delta\in[0,1)^{d}$. Given a lattice point set $\mathcal{P}_{N}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N-1}\}$ and drawing $\boldsymbol{\Delta}$ from a uniform distribution over $0,1)^{d}$, the corresponding randomly shifted lattice rule applied to a function $g$ is then given by
$Q_{N,d,\boldsymbol{\Delta}}(g):=\frac{1}{N}\sum_{k=0}^{N-1}g\left(\left\{\boldsymbol{x}_{k}+\boldsymbol{\Delta}\right\}\right),$
i.e., all points of $\mathcal{P}_{N}$ are shifted modulo one by the same $\boldsymbol{\Delta}$. If we would like to emphasize the role of the generating vector $\boldsymbol{z}$ of $\mathcal{P}_{N}$, we also write $Q_{N,d,\boldsymbol{z},\boldsymbol{\Delta}}$ in the following.
In the subsequent section, we will give our median algorithm based on lattice rules and analyze the corresponding approximation error.
3 The median algorithm and its error
As outlined in the introduction, we now turn to the definition of the median lattice-based algorithm, which will be used for $L_{2}$-approximation in the weighted Korobov space $\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$. Throughout this section, we assume downward-closed weights $\boldsymbol{\gamma}$, that is, $0<\gamma_{\mathfrak{v}}\leq\gamma_{\mathfrak{w}}\leq 1$ whenever $\mathfrak{w}\subseteq\mathfrak{v}\subseteq\{1{:}d\}$, and a smoothness parameter $\alpha>1/2$. The parameters $\boldsymbol{\gamma}$ and $\alpha$ are assumed to be fixed with these properties but otherwise arbitrary.
In this section, we first introduce the universal median lattice-based algorithm. We then state the main result of this paper, Theorem 3.8, which provides a probabilistic error bound for the algorithm. The proof of this result relies on the error decomposition given in ($\clubsuit$3.2), while the detailed analysis of the individual terms in the decomposition is deferred to the next section. Finally, we discuss the tractability properties of the median lattice-based algorithm.
3.1 The universal median lattice-based algorithm
For a real number $L\geq 0$, we define
$\mathcal{A}_{d}(L):=\left\{\boldsymbol{h}\in\mathbb{Z}^{d}:\prod_{j\in\operatorname{supp}(\boldsymbol{h})}\left|h_{j}\right|<L\right\},$
and
$\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L):=\left\{\boldsymbol{h}\in\mathbb{Z}^{d}:r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})<L^{2\alpha}\right\}=\left\{\boldsymbol{h}\in\mathbb{Z}^{d}:r_{1,\boldsymbol{\gamma}^{1/(2\alpha)}}(\boldsymbol{h})<L\right\},$
where $\boldsymbol{\gamma}^{1/(2\alpha)}:=(\gamma_{\mathfrak{u}}^{1/(2\alpha)})_{\mathfrak{u}\subseteq\{1{:}d\}}$. Note that $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)\subseteq\mathcal{A}_{d}(L)$, since $\sup_{\mathfrak{u}\subseteq\{1{:}d\}}\gamma_{\mathfrak{u}}\leq 1$ and $2\alpha>1$.
For an odd positive integer $R$, we define the median of $R$ complex numbers $Z_{1},\dots,Z_{R}$ by
$\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}(Z_{r}):=\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}\Re(Z_{r})+\mathtt{i}\,\cdot\,\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}\Im(Z_{r}),$
where $\Re(Z_{r})$ and $\Im(Z_{r})$ denote the real and imaginary parts of $Z_{r}$, respectively.
In the following, we also assume that $N$ is an odd prime number. Although this requirement is not necessary for all results presented below, it is convenient to make this assumption from the outset, as some statements do rely on it.
Algorithm 3.1.
For given $R,d\in\mathbb{N}$ and an odd prime $N$, perform the following steps:
  1. 1.
    For $r=1,\dots,R$, do:
    1. (a)
      Randomly draw $\boldsymbol{z}_{r}$ from the uniform distribution over $\{1{:}(N-1)\}^{d}$.
    2. (b)
      Randomly draw $\boldsymbol{\Delta}_{r}$ from the uniform distribution over $0,1)^{d}$.
    3. (c)
      For $\boldsymbol{h}\in\mathbb{Z}^{d}$, define
      $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}):=\frac{1}{N}\sum_{k=0}^{N-1}f\left(\left\{\frac{k\boldsymbol{z}_{r}}{N}+\boldsymbol{\Delta}_{r}\right\}\right)\exp\left(-2\pi\mathtt{i}\,\boldsymbol{h}\cdot\left(\frac{k\boldsymbol{z}_{r}}{N}+\boldsymbol{\Delta}_{r}\right)\right).$
  2. 2.
    Order $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ such that the medians of $|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ over $r\in\{1{:}R\}$ are in descending order, i.e.,
    $\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}_{1})|^{2}\geq\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}_{2})|^{2}\geq\cdots$
    and set $K_{N}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\}$.
  3. 3.
    Define the median estimator
    $\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h}):=\operatorname*{median}_{\begin{subarray}{c}r\in\{1{:}R\}\end{subarray}}\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}),$
    and the corresponding approximation
    $(A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f))(\boldsymbol{x}):=\sum_{\boldsymbol{h}\in K_{N}}\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})\exp(2\pi\mathtt{i}\,\boldsymbol{h}\cdot\boldsymbol{x}),$
    where $\boldsymbol{z}_{\{1{:}R\}}$ and $\boldsymbol{\Delta}_{\{1{:}R\}}$ denote the collections of $\boldsymbol{z}_{r}$ and $\boldsymbol{\Delta}_{r}$, $r=1,\dots,R$, respectively.
Remark 3.2.
The key difference from the algorithm proposed in [22] is that the present method does not require prior specification of the smoothness parameter $\alpha$ or the weights $\boldsymbol{\gamma}$. By considering the enlarged hyperbolic cross index set $\mathcal{A}_{d}(N/2)$ and sorting the estimated Fourier coefficients in descending order, the algorithm automatically adapts to the underlying smoothness and weights of the function.
Remark 3.3.
The total number of function evaluations is $M=RN$. Since we compute $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$ for all $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, an additional computational effort of order $O(|\mathcal{A}_{d}(N/2)|)$ is required. It is known that $|\mathcal{A}_{d}(N/2)|\asymp N(\log N)^{d-1}$ (see, for instance, [6, Chapter 2]), so that the associated computational cost can become significant for large $d$. Such an exponential dependence of the work on the dimension $d$ does not occur in the previous method studied in [22]. This highlights a trade-off between achieving universality and controlling the computational budget.
3.2 Error analysis: the main result
In this subsection, we first establish a few basic lemmas and then state the main result of this paper, Theorem 3.8, which provides a probabilistic bound on the $L_{2}$-approximation error. This result is based on the error decomposition given in ($\clubsuit$3.2). The detailed derivation of the bounds for the individual terms in this decomposition is deferred to the next section.
We first estimate the cardinality of $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)$. The following result was shown in [5].
Lemma 3.4.
Let $\boldsymbol{\gamma}=(\gamma_{\mathfrak{u}})_{\mathfrak{u}\subseteq\{1{:}d\}}$ be arbitrary positive weights. Then, for $\alpha>1/2$ and $\lambda>1/(2\alpha)$,
$\left|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)\right|\leq T_{\alpha,\lambda}(\boldsymbol{\gamma})\,L^{2\alpha\lambda},$
where
$T_{\alpha,\lambda}(\boldsymbol{\gamma}):=\sum_{\mathfrak{u}\subseteq\{1:d\}}\gamma_{\mathfrak{u}}^{\lambda}\,(2\zeta(2\alpha\lambda))^{\left|\mathfrak{u}\right|},$
and $\zeta(q):=\sum_{n=1}^{\infty}n^{-q}$ denotes the Riemann zeta function.
Next, for an odd prime $N$, we define
$N_{*}:=\sup\left\{L\geq 0:|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|\leq\frac{N-1}{16}\right\}.$ (2)
Since $|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(0)|=0$, the above set is nonempty and thus $N_{*}>0$. By definition, we then have
$|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})|\leq\frac{N-1}{16}.$ (3)
Note that $N_{*}<N/2$ because there are already $N$ integers between N/2-N/2 and $N/2$. We also have the following lower bound on $N_{*}$:
Lemma 3.5.
For an odd prime $N$, $\alpha>1/2$, $\lambda>1/(2\alpha)$, and positive weights $\boldsymbol{\gamma}=(\gamma_{\mathfrak{u}})_{\mathfrak{u}\subseteq\{1{:}d\}}$,
$N_{*}\geq\left(\frac{N-1}{16\,T_{\alpha,\lambda}(\boldsymbol{\gamma})}\right)^{1/(2\alpha\lambda)}.$
Proof.
Let $L>N_{*}$. By Lemma 3.4 and the definition of $N_{*}$, we have
$\frac{N-1}{16}<|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|\leq T_{\alpha,\lambda}(\boldsymbol{\gamma})\,L^{2\alpha\lambda}.$
Taking the limit $L\downarrow N_{*}$ then yields the desired bound. ∎
Intuitively, $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})$ contains all the large Fourier coefficients that we want our algorithm to estimate and include in $K_{N}$. The next lemma provides a bound on the error arising from leaving out frequencies that are not in $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})$.
Lemma 3.6.
For $L>0$ and $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$,
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)}|\widehat{f}(\boldsymbol{h})|^{2}\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{L^{2\alpha}}.$
Proof.
We have
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)}|\widehat{f}(\boldsymbol{h})|^{2}\leq\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)}|\widehat{f}(\boldsymbol{h})|^{2}\frac{r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})}{L^{2\alpha}}\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{L^{2\alpha}}.\qed$
In the following, we write
$N\cdot\mathbb{Z}^{d}:=\{\boldsymbol{\ell}\in\mathbb{Z}^{d}:\ell_{j}\equiv 0{\;(\mathrm{mod}\;N)}\text{ for all }j\in\{1{:}d\}\}.$
The next result is an immediate corollary of Lemma 3.6.
Corollary 3.7.
For an odd prime $N$ and $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$,
$\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(N/2)}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}.$
Proof.
Since $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ implies $\boldsymbol{h}\in(-N/2,N/2)^{d}$ for an odd prime $N$, the shifted lattices $\boldsymbol{h}+N\cdot\mathbb{Z}^{d}$ are mutually disjoint for different $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$. Moreover, for any $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and $\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$, we have $\boldsymbol{h}+\boldsymbol{\ell}\notin\mathcal{A}_{d}(N/2)$. Hence, by Lemma 3.6,
$\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(N/2)}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\leq\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d}(N/2)}|\widehat{f}(\boldsymbol{h})|^{2}\leq\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N/2)}|\widehat{f}(\boldsymbol{h})|^{2}\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{(N/2)^{2\alpha}}.$
The result then follows from the fact that $N_{*}<N/2$. ∎
To derive a probabilistic error bound for the algorithm $A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}$, we introduce some additional notation and an auxiliary set of indices. For the set $K_{N}$ generated by our algorithm, define
$f_{K_{N}}:=\sum_{\boldsymbol{h}\in K_{N}}\widehat{f}(\boldsymbol{h})\exp(2\pi\mathtt{i}\boldsymbol{h}\cdot\boldsymbol{x}).$
Moreover, consider an alternative approximation to $f$ using an oracle with access to all Fourier coefficients of $f$. For $\boldsymbol{h}^{\prime}\in\mathcal{A}_{d}(N/2)$, define
$F(\boldsymbol{h}^{\prime}):=\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})|^{2}.$ (4)
Ordering the elements $\boldsymbol{h}^{\prime}$ of $\mathcal{A}_{d}(N/2)$ so that $F(\boldsymbol{h}^{\prime}_{1})\geq F(\boldsymbol{h}^{\prime}_{2})\geq\cdots$, we set $K^{\prime}_{N}=\{\boldsymbol{h}^{\prime}_{1},\dots,\boldsymbol{h}^{\prime}_{N}\}$.
The main idea of the remaining proof is to use the triangle inequality to obtain an upper bound on the squared error:
$\|f-A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f)\|^{2}_{L_{2}}$
$\leq\Big(\|f_{K_{N}}-A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f)\|_{L_{2}}+\|f-f_{K_{N}}\|_{L_{2}}\Big)^{2}$
$\leq 2\|f_{K_{N}}-A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f)\|^{2}_{L_{2}}+2\|f-f_{K_{N}}\|^{2}_{L_{2}}$
$=2\sum_{\boldsymbol{h}\in K_{N}}\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}+2\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus K_{N}}|\widehat{f}(\boldsymbol{h})|^{2}$
$\leq 2\sum_{\boldsymbol{h}\in K_{N}}\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}+2\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus K^{\prime}_{N}}|\widehat{f}(\boldsymbol{h})|^{2}+2\sum_{\boldsymbol{h}\in K^{\prime}_{N}\setminus K_{N}}|\widehat{f}(\boldsymbol{h})|^{2},$ ($\clubsuit$)
where we have used the orthonormality of the Fourier basis in the equality.
We will analyze the three summands in the last line separately in the next section. The corresponding probabilistic bounds are established in Theorems 4.5, 4.6, and 4.14, respectively. By combining these results, we arrive at the following main theorem.
Theorem 3.8.
Let $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, and let $N$ be an odd prime and $R$ an odd natural number, chosen sufficiently large such that $\varepsilon_{1}+\varepsilon_{2}\leq\varepsilon$ for some $\varepsilon\in(0,1)$, where $\varepsilon_{1}$ and $\varepsilon_{2}$ are defined in (12) and (22), respectively. Then, with probability at least $1-\varepsilon$, we have
$\|f-A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f)\|^{2}_{L_{2}}\leq\frac{32\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}(2N-1)}{N^{2\alpha}_{*}(N-1)}\left(\frac{991N-255}{120(2N-1)}+\max\!\left(960,\sqrt{240N^{4\alpha}_{*}\bar{r}_{\mathrm{sum}}}\right)\right),$
where $N_{*}$ and $\bar{r}_{\mathrm{sum}}$ are as defined in (2) and (17), respectively.
With the $(\epsilon,\delta)$-approximation framework (1) in mind, our randomized algorithm achieves an $(\epsilon,\delta)$-approximation with
$\epsilon^{2}=\frac{32(2N-1)}{N^{2\alpha}_{*}(N-1)}\left(\frac{991N-255}{120(2N-1)}+\max\!\left(960,\sqrt{240N^{4\alpha}_{*}\bar{r}_{\mathrm{sum}}}\right)\right),$
for $\delta\leq\varepsilon_{1}+\varepsilon_{2}\in(0,1)$.
Remark 3.9.
Given a failure probability tolerance $\varepsilon\in(0,1)$, the condition $\varepsilon_{1}+\varepsilon_{2}\leq\varepsilon$ holds, where $\varepsilon_{1}$ and $\varepsilon_{2}$ are given in (12) and (22), respectively, provided that
$N\left(\frac{1}{2}\right)^{R/2}+\frac{|\mathcal{A}_{d}(N/2)|}{2}\left(\frac{1}{2}\right)^{R/2}+\frac{N}{2}\left(\frac{3}{4}\right)^{R/2}\leq\varepsilon.$
This inequality is satisfied if
$N\left(\frac{3}{4}\right)^{R/2}\leq\frac{\varepsilon}{3}\quad\text{and}\quad|\mathcal{A}_{d}(N/2)|\left(\frac{1}{2}\right)^{R/2}\leq\varepsilon.$
The first condition holds whenever $R\geq 2\log(3N/\varepsilon)/(\log(4/3))$. Next, since $\mathcal{A}_{d}(N/2)$ is a special instance of $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)$, Lemma 3.4 with $L=N/2$, $\lambda=1/\alpha$, and $\gamma_{u}=1$ for all $u\subseteq\{1{:}d\}$ implies
$|\mathcal{A}_{d}(N/2)|\leq\frac{N^{2}}{4}\,(1+2\zeta(2))^{d}.$
Thus, the second condition holds if
$R\;\geq\;2\left(\frac{\log(3N^{2})-\log(4\varepsilon)}{\log(2)}+\frac{\log(1+2\zeta(2))}{\log(2)}\,d\right).$
We conclude that the error bound in Theorem 3.8 fails with probability at most $\varepsilon$ if
$R\geq 2\max\left(\frac{\log(3N/\varepsilon)}{\log(4/3)},\;\frac{\log(3N^{2})-\log(4\varepsilon)}{\log(2)}+\frac{\log(1+2\zeta(2))}{\log(2)}\,d\right).$ (5)
3.3 Tractability analysis
The probabilistic bound for the $L_{2}$-approximation error established in Theorem 3.8 still depends on the quantities $N_{*}$ and $\bar{r}_{\mathrm{sum}}$, and thus does not yet yield an explicit error bound that is independent of these. In this subsection, we analyze these quantities under suitable assumptions and derive bounds that either depend only polynomially on the dimension $d$ or are independent of $d$. This, in turn, allows us to investigate the tractability properties of the universal median lattice-based algorithm.
First, let us consider the quantity $\bar{r}_{\mathrm{sum}}$ defined in (17). By Lemma 3.4, for any $\lambda\in(1/(2\alpha),1)$ we obtain
$\int_{N^{\ast}}^{\infty}\frac{\left|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)\right|}{L^{4\alpha+1}}\,\mathrm{d}L\leq T_{\alpha,\lambda}(\boldsymbol{\gamma})\int_{N^{\ast}}^{\infty}\frac{1}{L^{4\alpha+1-2\alpha\lambda}}\,\mathrm{d}L=\frac{T_{\alpha,\lambda}(\boldsymbol{\gamma})}{2\alpha(2-\lambda)}\,\frac{1}{N_{*}^{4\alpha-2\alpha\lambda}}.$
This implies that
$\bar{r}_{\mathrm{sum}}\leq\frac{32}{2-\lambda}\,\frac{T_{\alpha,\lambda}(\boldsymbol{\gamma})}{(N-1)N_{*}^{4\alpha-2\alpha\lambda}}+\frac{16}{N^{4\alpha}}\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1:d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|}.$
Moreover, since $N_{*}<N/2$ always holds, we further obtain
$N^{4\alpha}_{*}\bar{r}_{\mathrm{sum}}\leq\frac{32}{2-\lambda}\,\frac{N^{2\alpha\lambda}T_{\alpha,\lambda}(\boldsymbol{\gamma})}{2^{2\alpha\lambda}(N-1)}+\frac{16}{2^{4\alpha}}\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1:d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|}.$
Combining the above estimate with Lemma 3.5 and Theorem 3.8, we obtain the following result.
Corollary 3.10.
For any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$ and $\varepsilon\in(0,1)$, if the odd prime $N$ and the odd natural number $R$ satisfy condition (5), then with probability at least $1-\varepsilon$, we have
$\|f-A^{\mathrm{rand}}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(f)\|^{2}_{L_{2}}$
$\leq\left\|f\right\|^{2}_{d,\alpha,\boldsymbol{\gamma}}\frac{C_{1}}{N^{2\alpha}_{*}}\left(1+C_{2}\sqrt{N^{4\alpha}_{*}\bar{r}_{\mathrm{sum}}}\right)$
$\leq\left\|f\right\|^{2}_{d,\alpha,\boldsymbol{\gamma}}\frac{C_{1,\alpha,\lambda}(T_{\alpha,\lambda}(\boldsymbol{\gamma}))^{1/\lambda}}{N^{1/\lambda}}\left(1+C_{2,\alpha,\lambda}\sqrt{N^{2\alpha\lambda-1}T_{\alpha,\lambda}(\boldsymbol{\gamma})+\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1:d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|}}\right),$ (6)
for any $\lambda\in(1/(2\alpha),1)$, where $C_{1},C_{2},C_{1,\alpha,\lambda},C_{2,\alpha,\lambda}$ are constants independent of $N$, $R$, and $d$, but possibly depending on $\alpha$ and $\lambda$, as indicated by their subscripts.
From Corollary 3.10, we can draw several important conclusions regarding the median lattice-based algorithm. First, by choosing $\lambda$ arbitrarily close to $1/(2\alpha)$, we can achieve, for any given $f$, a convergence rate for the $L_{2}$-approximation error that is arbitrarily close to the optimal rate $N^{-\alpha}$, with high probability. Since our algorithm does not require any prior knowledge of the function’s smoothness or the weight parameters, this near-optimal convergence rate applies universally for any smoothness $\alpha>1/2$ and any downward-closed weights $\boldsymbol{\gamma}$.
Next, for $q\geq 0$, define
$B_{\lambda,q}:=\sup_{d\in\mathbb{N}}\left(\frac{1}{d^{q}}\,T_{\alpha,\lambda}(\boldsymbol{\gamma})\right)\quad\text{and}\quad\widetilde{B}_{q}:=\sup_{d\in\mathbb{N}}\left(\frac{1}{d^{q}}\,\sum_{\mathfrak{u}\subseteq\{1:d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|}\right).$
If, for some $\lambda\in(1/(2\alpha),1)$ and some $q>0$, we have $B_{\lambda,q}<\infty$ and $\widetilde{B}_{q}<\infty$, then the error bound in (6) depends at most polynomially on $d$, and an error of at most $\varepsilon>0$ can be achieved with $N$ depending at most polynomially on $d$ and $\varepsilon$, with high probability. Moreover, the number $R$ of independent rank-1 lattices required for this result depends only logarithmically on $N$ and linearly on $d$ by (5). This situation is referred to as polynomial tractability in the context of Information-Based Complexity (see, e.g., [21]).
In the case of product weights, where $\gamma_{\mathfrak{u}}=\prod_{j\in\mathfrak{u}}\gamma_{j}$ with $\gamma_{j}\leq 1$, we have
$T_{\alpha,\lambda}(\boldsymbol{\gamma})=\sum_{\mathfrak{u}\subseteq\{1:d\}}\prod_{j\in u}\left(2\gamma_{j}^{\lambda}\zeta(2\alpha\lambda)\right)=\prod_{j=1}^{d}\left(1+2\gamma_{j}^{\lambda}\zeta(2\alpha\lambda)\right).$
Assume that
$A_{\lambda}:=\limsup_{d\to\infty}\frac{1}{\log d}\sum_{j=1}^{d}\gamma_{j}^{\lambda}<\infty,$ (7)
for some $\lambda\in(1/(2\alpha),1)$. This assumption implies that, for any fixed $\delta>0$, there exists $d_{\delta}$ such that
$\frac{1}{\log d}\sum_{j=1}^{d}\gamma_{j}^{\lambda}<A_{\lambda}+\delta,\qquad\text{for all $d\geq d_{\delta}$.}$
Applying the elementary inequality $\exp(x)\geq x+1$, we then obtain
$\frac{1}{d^{q}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{\lambda}\zeta(2\alpha\lambda)\right)$ $\leq\frac{1}{d^{q}}\exp\left(2\zeta(2\alpha\lambda)\sum_{j=1}^{d}\gamma_{j}^{\lambda}\right)$
$\leq\frac{1}{d^{q}}\exp\left(2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)\log d)\right)=d^{2(A_{\lambda}+\delta)\zeta(2\alpha\lambda)-q}.$
Hence, $B_{\lambda,q}<\infty$ for any $q>2A_{\lambda}\zeta(2\alpha\lambda)$. Similarly, if (7) holds with $\lambda=2$, then $\widetilde{B}_{q}<\infty$ for any $q>2A_{2}\zeta(4\alpha)$.
If $B_{\lambda,q}<\infty$ and $\widetilde{B}_{q}<\infty$ even hold for $q=0$, then the error bound becomes independent of $d$. In the case of product weights, these conditions simplify to
$\sum_{j=1}^{\infty}\gamma_{j}^{\lambda}<\infty\quad\text{and}\quad\sum_{j=1}^{\infty}\gamma_{j}^{2}<\infty,$
respectively. Unfortunately, the required number $R$ still depends linearly on $d$ (cf. (5)), which prevents us from claiming strong polynomial tractability for our approximation algorithm. This limitation appears to be a consequence of the universality of the algorithm. Nevertheless, the linear dependence on $d$ may be acceptable for applications with moderate dimension. If strong tractability is desired, one could instead use the algorithm from [22], which achieves the same near-optimal convergence order but does not maintain universality.
4 Error analysis: bounds on the error decomposition
In this section, we provide a detailed analysis of the three summands in the error decomposition ($\clubsuit$3.2), which underlies the probabilistic $L_{2}$-error bounds given in Theorem 3.8 and Corollary 3.10. Each term is treated in a separate subsection, where we establish probabilistic bounds that, when combined, recover the main result presented in Section 3.
4.1 The first term of ($\clubsuit$3.2)
Our proof in this part is adapted from that in [22]. First, we show that, for most random choices of $\boldsymbol{z}_{r}$, the generated lattices are “good” for estimating $\widehat{f}(\boldsymbol{h})$.
Lemma 4.1.
For any $r\in\{1{:}R\}$, define
$P_{N,\boldsymbol{z}_{r}}^{\perp}:=\{\boldsymbol{\ell}\in\mathbb{Z}^{d}:\boldsymbol{z}_{r}^{\top}\boldsymbol{\ell}\equiv 0{\;(\mathrm{mod}\;N)}\}.$ (8)
For any $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, define
$\mathcal{B}_{\boldsymbol{h}}:=\left\{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d}):r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})<N_{*}^{2\alpha}\right\},$
where $N_{*}$ is as defined in (2). Then
$\Pr(P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset)\leq\frac{1}{16}.$
Proof.
By (3), we have
$|\mathcal{B}_{\boldsymbol{h}}|$ $=\Big|\big\{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d}):r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})<N_{*}^{2\alpha}\big\}\Big|$
$\leq\Big|\big\{\boldsymbol{\ell}\in\mathbb{Z}^{d}:r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{\ell})<N_{*}^{2\alpha}\big\}\Big|\leq\frac{N-1}{16}.$
For each $\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d})$,
$\Pr(\boldsymbol{z}_{r}^{\top}\boldsymbol{\ell}\equiv 0{\;(\mathrm{mod}\;N)})=\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}N-1\}^{d}}\mathbf{1}\{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\}\leq\frac{1}{N-1}.$ (9)
It then follows from a union bound that
$\Pr(P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset)\leq\frac{|\mathcal{B}_{\boldsymbol{h}}|}{N-1}\leq\frac{1}{16}.\qed$
The next lemma bounds the typical error in replacing $\widehat{f}(\boldsymbol{h})$ with $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$.
Lemma 4.2.
Let $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$. For $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, define
$\delta(\boldsymbol{h}):=4\left(\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\right)^{1/2}.$
Then, for any $r\in\{1{:}R\}$,
$\Pr\left(|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}>\delta(\boldsymbol{h})^{2}\right)\leq\frac{1}{8}.$
Proof.
By Lemma 4.1 and Markov’s inequality, it holds that
$\Pr\left(|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}>\delta(\boldsymbol{h})^{2}\right)$
$\leq\Pr(P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset)+\Pr\left(\{|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}>\delta(\boldsymbol{h})^{2}\}\cap\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}\right)$
$=\Pr(P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset)+\Pr\left(\delta(\boldsymbol{h})^{-2}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}>1\right)$
$\leq\frac{1}{16}+\frac{1}{\delta(\boldsymbol{h})^{2}}\mathbb{E}\left[|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}\right].$ (10)
As outlined in [3, Proof of Lemma 4.1],
$\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}\,d\boldsymbol{\Delta}_{r}=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}.$ (11)
Taking expectation over $\boldsymbol{z}_{r}$ and using independence of $\boldsymbol{\Delta}_{r}$ and $\boldsymbol{z}_{r}$, we obtain
$\mathbb{E}\left[|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}\right]$
$=\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}$
$\leq\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}\left(\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus(N\cdot\mathbb{Z}^{d})}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\frac{r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}{N_{*}^{2\alpha}}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\right)$
$=\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d})}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\frac{r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}{N_{*}^{2\alpha}}\left(\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}\boldsymbol{1}\{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\}\right)+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}$
$\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2},$
where we have used (9). Substituting this bound into (4.1) yields the desired probability bound. ∎
The next lemma illustrates how the median trick reduces the failure probability. It comes from [16, Proposition 2.2], which is a minor modification of [20, Proposition 2.1].
Lemma 4.3.
Let $R$ be an odd natural number, and let $X_{1},X_{2},\dots,X_{R}$ be $R$ independent and identically distributed real random variables. Suppose that for some interval $I\subseteq\mathbb{R}$ and $p\in(0,1/2)$ we have
$\Pr\left(X_{1}\notin I\right)\leq p.$
Then
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}(X_{r})\notin I\right)\leq\frac{1}{2}\,(4p(1-p))^{R/2}.$
The following result is a direct consequence of Lemmas 4.2 and 4.3.
Lemma 4.4.
Suppose that $R$ is an odd integer and $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$. Then
$\Pr\left(\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}>2\,\delta(\boldsymbol{h})^{2}\right)\leq 2^{-R/2},$
where $\delta(\boldsymbol{h})$ is as defined in Lemma 4.2.
Proof.
By Lemma 4.2, for each $r\in\{1{:}R\}$ we have
$\Pr\left(|\Re\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})-\Re\widehat{f}(\boldsymbol{h})|^{2}>\delta(\boldsymbol{h})^{2}\right)\leq\frac{1}{8}.$
Applying Lemma 4.3, it follows that
$\Pr\left(\left|\Re\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\Re\widehat{f}(\boldsymbol{h})\right|^{2}>\delta(\boldsymbol{h})^{2}\right)\leq\frac{1}{2}\left(4\cdot\frac{1}{8}\cdot\frac{7}{8}\right)^{R/2}\leq\frac{1}{2}2^{-R/2}.$
A similar bound holds for the imaginary part. The claim then follows by applying a union bound. ∎
We now show a bound on the first summand of ($\clubsuit$3.2).
Theorem 4.5.
Let $N$ be an odd prime and let $R$ be an odd natural number satisfying
$\varepsilon_{1}:=N\,2^{-R/2}<1.$ (12)
Then, for any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, with probability at least $1-\varepsilon_{1}$ we have
$\sum_{\boldsymbol{h}\in K_{N}}\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}\leq 32\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}\;\frac{2N-1}{N-1}.$
Proof.
By Lemma 4.4,
$\Pr\left(\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}>2\delta(\boldsymbol{h})^{2}\text{ for at least one }\boldsymbol{h}\in K_{N}\right)\leq|K_{N}|\,\,2^{-R/2}=\varepsilon_{1}.$
Hence, with probability at least $1-\varepsilon_{1}$, we have $\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}\leq 2\delta(\boldsymbol{h})^{2}$ for all $\boldsymbol{h}\in K_{N}$. It then follows that
$\sum_{\boldsymbol{h}\in K_{N}}\left|\widehat{f}_{N,\boldsymbol{z}_{\{1{:}R\}},\boldsymbol{\Delta}_{\{1{:}R\}}}(\boldsymbol{h})-\widehat{f}(\boldsymbol{h})\right|^{2}$ $\leq 2\sum_{\boldsymbol{h}\in K_{N}}\delta(\boldsymbol{h})^{2}$
$\leq 32\left(\frac{|K_{N}|}{N-1}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}+\sum_{\boldsymbol{h}\in K_{N}}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\right)$
$\leq 32\left(\frac{N}{N-1}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}+\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}\right),$
where the last inequality uses Corollary 3.7. ∎
4.2 The second term of ($\clubsuit$3.2)
Next, we turn to the second term in the error decomposition ($\clubsuit$3.2). Recall that $F(\boldsymbol{h}^{\prime})$ is defined in (4) and that the set $K^{\prime}_{N}=\{\boldsymbol{h}^{\prime}_{1},\dots,\boldsymbol{h}^{\prime}_{N}\}\subset\mathcal{A}_{d}(N/2)$ consists of the indices corresponding to the $N$ largest values of $F(\boldsymbol{h}^{\prime})$. Consequently, the contribution of the Fourier coefficients outside $K^{\prime}_{N}$ can be suitably bounded, as proven in the following theorem.
Theorem 4.6.
For any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, we have
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus K^{\prime}_{N}}|\widehat{f}(\boldsymbol{h})|^{2}\leq\frac{2\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N^{2\alpha}_{*}},$
where $N_{*}$ is as defined in (2).
Proof.
Since $\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})\subseteq\mathcal{A}_{d}(N/2)$ and $|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})|\leq N$, it follows that
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})}|\widehat{f}(\boldsymbol{h})|^{2}=$ $\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h})|^{2}-\sum_{\boldsymbol{h}\in\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})}|\widehat{f}(\boldsymbol{h})|^{2}$
$\geq$ $\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h})|^{2}-\sum_{\boldsymbol{h}\in\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})}F(\boldsymbol{h})$
$\geq$ $\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h})|^{2}-\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}F(\boldsymbol{h}^{\prime})$
$=$ $\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h})|^{2}-\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})|^{2}$
$=$ $\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus K^{\prime}_{N}}|\widehat{f}(\boldsymbol{h})|^{2}-\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})|^{2}.$
Therefore,
$\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus K^{\prime}_{N}}|\widehat{f}(\boldsymbol{h})|^{2}\leq\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}\setminus\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(N_{*})}|\widehat{f}(\boldsymbol{h})|^{2}+\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})|^{2}.$
The desired bound follows by applying Lemma 3.6 and Corollary 3.7. ∎
4.3 The third term of ($\clubsuit$3.2)
Finally, we consider the third term in the error decomposition ($\clubsuit$3.2). In contrast to the first two terms, bounding this term requires a more detailed analysis. The key idea is to show that, with high probability, all $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}$ corresponding to large values of $|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}$ are contained in $K_{N}$.
To this end, we first establish several lemmas to study the typical size of $|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$, focusing in particular on its variance with respect to the random shift $\boldsymbol{\Delta}_{r}$ conditioned on a fixed $\boldsymbol{z}_{r}$.
Lemma 4.7.
For any $r\in\{1{:}R\}$ and $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, we have
$\int_{[0,1)^{d}}\bigl|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\bigr|^{2}\,\mathrm{d}\boldsymbol{\Delta}_{r}$ $=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2},$ (13)
$\int_{[0,1)^{d}}\bigl|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\bigr|^{4}\,\mathrm{d}\boldsymbol{\Delta}_{r}$ $=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\,\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2},$ (14)
where $P_{N,\boldsymbol{z}_{r}}^{\perp}$ denotes the dual lattice of $P_{N,\boldsymbol{z}_{r}}$, as defined in (8).
Proof.
The first equality follows similarly to (11). For the second equality, we start from [3, Proof of Lemma 4.1]:
$\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})\,\exp(2\pi\mathtt{i}\boldsymbol{\ell}\cdot\boldsymbol{\Delta}_{r}).$
It follows that
$|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{4}=\left|\sum_{\boldsymbol{\ell},\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\exp(2\pi\mathtt{i}(\boldsymbol{\ell}+\boldsymbol{\ell}^{\prime})\cdot\boldsymbol{\Delta}_{r})\right|^{2}.$
Setting $\boldsymbol{L}=\boldsymbol{\ell}+\boldsymbol{\ell}^{\prime}$ and defining
$\widehat{f}_{*}(\boldsymbol{L}):=\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{L}-\boldsymbol{\ell}^{\prime})\,\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime}),$
we can rewrite the double sum as
$\sum_{\boldsymbol{\ell},\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})\,\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\,e^{2\pi i(\boldsymbol{\ell}+\boldsymbol{\ell}^{\prime})\cdot\boldsymbol{\Delta}_{r}}=\sum_{\boldsymbol{L}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}_{*}(\boldsymbol{L})\exp(2\pi\mathtt{i}\boldsymbol{L}\cdot\boldsymbol{\Delta}_{r}).$
Finally, integrating over $\boldsymbol{\Delta}_{r}\in[0,1)^{d}$ and using the orthonormality of the Fourier basis gives
$\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{4}\,\mathrm{d}\boldsymbol{\Delta}_{r}=\int_{[0,1)^{d}}\left|\sum_{\boldsymbol{L}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}_{*}(\boldsymbol{L})\exp(2\pi\mathtt{i}\boldsymbol{L}\cdot\boldsymbol{\Delta}_{r})\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}_{r}=\sum_{\boldsymbol{L}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}|\widehat{f}_{*}(\boldsymbol{L})|^{2}.\qed$
Lemma 4.8.
For any $r\in\{1{:}R\}$ and $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, define
$r_{\min}(\boldsymbol{h})$ $:=\min_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}),$
$r_{\mathrm{sum}}(\boldsymbol{h})$ $:=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}\frac{1}{r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})},$
$S(\boldsymbol{h})$ $:=\left(\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{1/2}.$
Then we have
$\sigma^{2}(\boldsymbol{h}):=\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{4}\,\mathrm{d}\boldsymbol{\Delta}_{r}-\left(\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\,\mathrm{d}\boldsymbol{\Delta}_{r}\right)^{2}\leq\frac{8|\widehat{f}(\boldsymbol{h})|^{2}}{r_{\min}(\boldsymbol{h})}S(\boldsymbol{h})^{2}+2r_{\mathrm{sum}}(\boldsymbol{h})S(\boldsymbol{h})^{4}.$
Proof.
We first bound the $\boldsymbol{\ell}=\boldsymbol{0}$ summand in (14):
$\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}-\boldsymbol{\ell}^{\prime})\right|\leq\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\frac{|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}+|\widehat{f}(\boldsymbol{h}-\boldsymbol{\ell}^{\prime})|^{2}}{2}=\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}.$
Then, by (13) and (14),
$\sigma^{2}(\boldsymbol{h})\leq\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}.$ (15)
For any $\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}$, we can decompose
$\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})=2\widehat{f}(\boldsymbol{h})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})+\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime}).$
By the Cauchy–Schwarz inequality,
$\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}$
$=\left|2\widehat{f}(\boldsymbol{h})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})+\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}$
$\leq 8|\widehat{f}(\boldsymbol{h})|^{2}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}+2\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}.$ (16)
The second term can be further bounded by
$\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}$
$\leq\left(\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\frac{1}{r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})}\right)$
$\qquad\times\left(\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right).$
Since we have
$\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\frac{1}{r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})}$
$\leq\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\frac{1}{2r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})}+\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\frac{1}{2r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})}\leq r_{\mathrm{sum}}(\boldsymbol{h}),$
for any $\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}$, we obtain
$\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}$
$\leq r_{\mathrm{sum}}(\boldsymbol{h})\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}\,\,\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})$
$=r_{\mathrm{sum}}(\boldsymbol{h})\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0},\boldsymbol{\ell}^{\prime}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})$
$\leq r_{\mathrm{sum}}(\boldsymbol{h})\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})S(\boldsymbol{h})^{2}=r_{\mathrm{sum}}(\boldsymbol{h})S(\boldsymbol{h})^{4}.$
It follows from (4.3) and the above bound that
$\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}\left|\sum_{\boldsymbol{\ell}^{\prime}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}^{\prime})\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell}-\boldsymbol{\ell}^{\prime})\right|^{2}$ $\leq 8|\widehat{f}(\boldsymbol{h})|^{2}\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}+2r_{\mathrm{sum}}(\boldsymbol{h})S(\boldsymbol{h})^{4}$
$\leq\frac{8|\widehat{f}(\boldsymbol{h})|^{2}}{r_{\min}(\boldsymbol{h})}S(\boldsymbol{h})^{2}+2r_{\mathrm{sum}}(\boldsymbol{h})S(\boldsymbol{h})^{4},$
which completes the proof because of (15). ∎
Lemma 4.9.
Let $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$. Define
$\bar{S}(\boldsymbol{h}):=4\left(\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\right)^{1/2},$
and
$\bar{r}_{\mathrm{sum}}=16\left(\frac{4\alpha}{N-1}\int_{N_{*}}^{\infty}\frac{|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|}{L^{4\alpha+1}}\,\mathrm{d}L+\frac{1}{N^{4\alpha}}\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|}\right).$ (17)
Then, for any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, the following bounds hold:
$\Pr\left(S(\boldsymbol{h})^{2}\geq\bar{S}(\boldsymbol{h})^{2}\right)\leq\frac{1}{16}$ (18)
and
$\Pr\left(\{r_{\min}(\boldsymbol{h})<N^{2\alpha}_{*}\}\cup\{r_{\mathrm{sum}}(\boldsymbol{h})\geq\bar{r}_{\mathrm{sum}}\}\right)\leq\frac{1}{8},$
where $N_{*}$ and $r_{\mathrm{sum}}(\boldsymbol{h})$ are as defined in (2) and Lemma 4.8, respectively.
Proof.
First, we compute
$\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}S(\boldsymbol{h})^{2}$
$=\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})$
$=\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d})}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\left(\frac{1}{(N-1)^{d}}\sum_{\boldsymbol{z}_{r}\in\{1{:}(N-1)\}^{d}}\boldsymbol{1}\{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\}\right)$
$\qquad+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})$
$\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}),$
where we have used the bound (9). Then, the first statement (18) follows from Markov’s inequality.
Next, since $N_{*}<N/2$ and $r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})>(N/2)^{2\alpha}$ for $\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$, we have $r_{\min}(\boldsymbol{h})<N_{*}^{2\alpha}$ if and only if $P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset$. Hence, $\Pr(r_{\min}(\boldsymbol{h})<N_{*}^{2\alpha})\leq 1/16$ by Lemma 4.1.
Finally, to bound $\Pr(r_{\mathrm{sum}}(\boldsymbol{h})\geq\bar{r}_{\mathrm{sum}})$, define
$R_{\boldsymbol{h}}(L):=|\{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus(N\cdot\mathbb{Z}^{d}):r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\leq L^{2\alpha}\}|.$
When $r_{\min}(\boldsymbol{h})\geq N_{*}^{2\alpha}$, integration by parts for the Riemann–Stieltjes integral yields
$\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}\setminus(N\cdot\mathbb{Z}^{d})}\frac{1}{r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$ $=\int_{N_{*}}^{\infty}L^{-4\alpha}\,\mathrm{d}R_{\boldsymbol{h}}(L)=R_{\boldsymbol{h}}(L)L^{-4\alpha}\bigg|_{N_{*}}^{\infty}-\int_{N_{*}}^{\infty}R_{\boldsymbol{h}}(L)\,\mathrm{d}L^{-4\alpha}$
$=4\alpha\int_{N_{*}}^{\infty}R_{\boldsymbol{h}}(L)L^{-4\alpha-1}\,\mathrm{d}L,$
where, in the last equality, we have used the fact that $R_{\boldsymbol{h}}(L)\leq|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|$ is dominated by $L^{4\alpha}$. Since it follows from (9) that
$\mathbb{E}[R_{\boldsymbol{h}}(L)]=\frac{1}{N-1}\,|\{\boldsymbol{\ell}\in\mathbb{Z}^{d}\setminus(N\cdot\mathbb{Z}^{d}):r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})\leq L^{2\alpha}\}|\leq\frac{|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|}{N-1},$
we obtain
$\mathbb{E}[r_{\mathrm{sum}}(\boldsymbol{h})\boldsymbol{1}\{r_{\min}(\boldsymbol{h})>N^{2\alpha}_{*}\}]\leq$ $\int_{N_{*}}^{\infty}\frac{|\mathcal{A}_{d,\alpha,\boldsymbol{\gamma}}(L)|}{N-1}4\alpha L^{-4\alpha-1}\,\mathrm{d}L+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{1}{r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}.$
It remains to bound the second term on the right-hand side. Since $\boldsymbol{h}=(h_{1},\dots,h_{d})\in\mathcal{A}_{d}(N/2)$, we have $|h_{j}|\leq N/2$ for all $j$. For $\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}$, write $\ell_{j}=k_{j}N$ with $k_{j}\in\mathbb{Z}$. If $k_{j}=0$ then $|\ell_{j}+h_{j}|\geq 0$, while if $k_{j}\neq 0$ then
$|\ell_{j}+h_{j}|\geq(|k_{j}|-1)N+\frac{N}{2},$
so in particular $\ell_{j}+h_{j}\neq 0$ when $k_{j}\neq 0$. Writing $(\boldsymbol{\ell}_{\mathfrak{u}},\boldsymbol{0}_{-\mathfrak{u}})$ for the vector with components $\ell_{j}$ if $j\in\mathfrak{u}$ and zero otherwise, the above observation implies that
$\mathrm{supp}(\boldsymbol{h}+(\boldsymbol{\ell}_{\mathfrak{u}},\boldsymbol{0}_{-\mathfrak{u}}))\supseteq\mathfrak{u},$
for any $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and $\boldsymbol{\ell}_{\mathfrak{u}}\in(N\cdot\mathbb{Z}\setminus\{0\})^{|\mathfrak{u}|}$. Thus, we have
$\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}\frac{1}{r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell})}$ $=\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\sum_{\boldsymbol{\ell}_{\mathfrak{u}}\in(N\cdot\mathbb{Z}\setminus\{\boldsymbol{0}\})^{\left|\mathfrak{u}\right|}}\frac{1}{r^{2}_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+(\boldsymbol{\ell}_{\mathfrak{u}},\boldsymbol{0}_{-\mathfrak{u}}))}$
$=\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\sum_{\boldsymbol{\ell}_{\mathfrak{u}}\in(N\cdot\mathbb{Z}\setminus\{\boldsymbol{0}\})^{\left|\mathfrak{u}\right|}}\gamma^{2}_{\mathrm{supp}(\boldsymbol{h}+(\boldsymbol{\ell}_{\mathfrak{u}},\boldsymbol{0}_{-\mathfrak{u}}))}\left|\boldsymbol{h}+(\boldsymbol{\ell}_{\mathfrak{u}},\boldsymbol{0}_{-\mathfrak{u}})\right|^{-4\alpha}$
$\leq\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\gamma^{2}_{\mathfrak{u}}\sum_{\boldsymbol{\ell}_{\mathfrak{u}}\in(N\cdot\mathbb{Z}\setminus\{\boldsymbol{0}\})^{\left|\mathfrak{u}\right|}}\prod_{j\in\mathfrak{u}}(h_{j}+\ell_{j})^{-4\alpha}$
$\leq\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\gamma^{2}_{\mathfrak{u}}\left(2\sum_{k=1}^{\infty}((k-1)N+N/2)^{-4\alpha}\right)^{\left|\mathfrak{u}\right|}$
$=\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\gamma^{2}_{\mathfrak{u}}\left(\frac{2^{1+4\alpha}}{N^{4\alpha}}\sum_{k=1}^{\infty}(2k-1)^{-4\alpha}\right)^{\left|\mathfrak{u}\right|}$
$\leq\frac{1}{N^{4\alpha}}\sum_{\emptyset\neq\mathfrak{u}\subseteq\{1{:}d\}}\gamma^{2}_{\mathfrak{u}}\left(2^{1+4\alpha}\zeta(4\alpha)\right)^{\left|\mathfrak{u}\right|},$
where we have used the assumption that $\gamma_{\mathfrak{v}}\leq\gamma_{\mathfrak{w}}$ whenever $\mathfrak{w}\subseteq\mathfrak{v}$. Combining the above expectation bound with Markov’s inequality, and recalling that $\Pr(r_{\min}(\boldsymbol{h})<N^{2\alpha}_{*})\leq 1/16$, we obtain the desired probability estimate. ∎
Now we are ready to show that, with high probability, every $\boldsymbol{h}\notin K^{\prime}_{N}$ has a small value of $\operatorname{median}_{r\in\{1{:}R\}}\lvert\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\rvert^{2}$, while every $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}$ with a large value of $\lvert\widehat{f}(\boldsymbol{h}^{\prime})\rvert^{2}$ also attains a correspondingly large $\operatorname{median}_{r\in\{1{:}R\}}\lvert\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}^{\prime})\rvert^{2}$. These two facts together imply that, except with small probability, the ordering of frequencies induced by the medians in Step 2 of Algorithm 3.1 correctly separates the significant Fourier coefficients (which should belong to $K_{N}$) from the insignificant ones (which should lie outside $K_{N}$); this is the key ingredient for controlling the third summand in ($\clubsuit$3.2).
Lemma 4.10.
For $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, let
$\delta^{\prime}(\boldsymbol{h})=4\left(\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}+F(\boldsymbol{h})\right)^{1/2},$
where $F(\boldsymbol{h})$ is defined in (4). Then
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}\left|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\right|^{2}>\delta^{\prime}(\boldsymbol{h})^{2}\right)\leq\frac{1}{2}2^{-R/2}.$
Proof.
By replacing (11) with (13), the same argument as described in the proof of Lemma 4.2 yields
$\mathbb{E}\left[|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}\right]\leq\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}+\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}=\frac{\delta^{\prime}(\boldsymbol{h})^{2}}{16}.$
Hence, using Lemma 4.1 together with Markov’s inequality (as in (4.1)), we obtain
$\Pr\left(|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}>\delta^{\prime}(\boldsymbol{h})^{2}\right)$
$\leq\Pr(P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}\neq\emptyset)+\frac{1}{\delta^{\prime}(\boldsymbol{h})^{2}}\mathbb{E}\left[|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\boldsymbol{1}\{P_{N,\boldsymbol{z}_{r}}^{\perp}\cap\mathcal{B}_{\boldsymbol{h}}=\emptyset\}\right]\leq\frac{1}{8}.$
Finally, applying Lemma 4.3 to the $R$ independent replicates gives the claimed probability bound. ∎
The next lemma provides an upper bound on $F(\boldsymbol{h}^{\prime}_{N})$, the smallest value of $F(\boldsymbol{h}^{\prime})$ among $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\subset\mathcal{A}_{d}(N/2)$, in terms of the norm $\|f\|_{d,\alpha,\boldsymbol{\gamma}}$.
Lemma 4.11.
For any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$,
$F(\boldsymbol{h}^{\prime}_{N})\leq\frac{16}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}.$
Proof.
First observe that for any $\boldsymbol{h}\in(-N/2,N/2)^{d}$,
$r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})=\min_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}+\boldsymbol{\ell}).$
Hence,
$\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}=\sum_{\boldsymbol{h}\in\mathbb{Z}^{d}}|\widehat{f}({\boldsymbol{h}})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})\geq\sum_{\boldsymbol{h}\in\mathcal{A}_{d}(N/2)}F(\boldsymbol{h})r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h})\geq F(\boldsymbol{h}^{\prime}_{N})\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}^{\prime}).$ (19)
Since $|K^{\prime}_{N}|=N$, while at most (N1)/16(N-1)/16 indices $\boldsymbol{h}^{\prime}\in\mathcal{A}_{d}(N/2)$ satisfy $r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}^{\prime})<N_{*}^{2\alpha}$ (by (3)), we obtain
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}^{\prime})\geq N_{*}^{2\alpha}\left(N-\frac{N-1}{16}\right)\geq\frac{15}{16}N_{*}^{2\alpha}(N-1).$
Inserting this estimate into (19) yields the desired inequality. ∎
Corollary 4.12.
For $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, define
$\delta^{\prime\prime}=4\left(\frac{31}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)}\right)^{1/2}.$
Then for all $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)\setminus K^{\prime}_{N}$,
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}\left|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})\right|^{2}>\delta^{\prime\prime 2}\right)\leq\frac{1}{2}\,2^{-R/2}.$
Proof.
If $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)\setminus K^{\prime}_{N}$, then by construction
$F(\boldsymbol{h})\leq F(\boldsymbol{h}^{\prime}_{N})\leq\frac{16}{15}\,\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}(N-1)},$
where the second inequality follows from Lemma 4.11. Substituting this bound into the definition of $\delta^{\prime}(\boldsymbol{h})$ in Lemma 4.2 gives $\delta^{\prime}(\boldsymbol{h})^{2}\leq\delta^{\prime\prime 2}$. The claimed probability estimate then follows immediately from Lemmas 4.2 and 4.3. ∎
Lemma 4.13.
For $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, if
$|\widehat{f}(\boldsymbol{h})|^{2}\geq\bar{S}(\boldsymbol{h})^{2}\max\left(960N^{-2\alpha}_{*},\sqrt{240\bar{r}_{\mathrm{sum}}}\right),$ (20)
where both $\bar{S}(\boldsymbol{h})$ and $\bar{r}_{\mathrm{sum}}$ are as defined in Lemma 4.9, then
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\frac{F(\boldsymbol{h})}{2}\right)\leq\frac{1}{2}\left(\frac{3}{4}\right)^{R/2}.$
Proof.
From (13), we have
$\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\,\mathrm{d}\boldsymbol{\Delta}_{r}=\sum_{\boldsymbol{\ell}\in P_{N,\boldsymbol{z}_{r}}^{\perp}}|\widehat{f}(\boldsymbol{h}+\boldsymbol{\ell})|^{2}\geq F(\boldsymbol{h})\geq|\widehat{f}(\boldsymbol{h})|^{2},$
which provides a lower bound on the expectation of $|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ when $\boldsymbol{z}_{r}$ is fixed and $\boldsymbol{\Delta}_{r}$ follows a uniform distribution over $0,1)^{d}$. Let $\sigma^{2}(\boldsymbol{h})$ denote the corresponding variance as in Lemma 4.8. Then, Cantelli’s inequality gives
$\int_{[0,1)^{d}}\boldsymbol{1}\left\{|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\frac{F(\boldsymbol{h})}{2}\right\}\,\mathrm{d}\boldsymbol{\Delta}_{r}$
$\leq\int_{[0,1)^{d}}\boldsymbol{1}\left\{|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\int_{[0,1)^{d}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\,\mathrm{d}\boldsymbol{\Delta}_{r}-\frac{F(\boldsymbol{h})}{2}\right\}\,\mathrm{d}\boldsymbol{\Delta}_{r}$
$\leq\frac{\sigma(\boldsymbol{h})^{2}}{(F(\boldsymbol{h})/2)^{2}+\sigma(\boldsymbol{h})^{2}}\leq\frac{1}{1+|\widehat{f}(\boldsymbol{h})|^{4}/(4\sigma(\boldsymbol{h})^{2})}.$ (21)
Under the “good lattice” event, i.e., when $r_{\min}(\boldsymbol{h})>N^{2\alpha}_{*}$, $r_{\mathrm{sum}}(\boldsymbol{h})<\bar{r}_{\mathrm{sum}}$, and $S(\boldsymbol{h})^{2}<\bar{S}(\boldsymbol{h})^{2}$, Lemma 4.8 gives
$\frac{\sigma^{2}(\boldsymbol{h})}{|\widehat{f}(\boldsymbol{h})|^{4}}\leq 8N^{-2\alpha}_{*}\frac{\bar{S}(\boldsymbol{h})^{2}}{|\widehat{f}(\boldsymbol{h})|^{2}}+2\bar{r}_{\mathrm{sum}}\frac{\bar{S}^{4}(\boldsymbol{h})}{|\widehat{f}(\boldsymbol{h})|^{4}}.$
By condition (20), the right-hand side is at most $1/60$, so the right-most side of (4.3) is bounded by $1/16$. Hence, by Lemma 4.9,
$\Pr\left(\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\frac{F(\boldsymbol{h})}{2}\right)$
$\leq\Pr\left(\{r_{\min}(\boldsymbol{h})\leq N^{2\alpha}_{*}\}\cup\{r_{\mathrm{sum}}(\boldsymbol{h})\geq\bar{r}_{\mathrm{sum}}\}\cup\{S(\boldsymbol{h})^{2}\geq\bar{S}(\boldsymbol{h})^{2}\}\right)$
$\qquad+\Pr(\{|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq F(\boldsymbol{h})/2\}\cap\{r_{\min}(\boldsymbol{h})>N^{2\alpha}_{*}\}\cap\{r_{\mathrm{sum}}(\boldsymbol{h})<\bar{r}_{\mathrm{sum}}\}\cap\{S(\boldsymbol{h})^{2}<\bar{S}(\boldsymbol{h})^{2}\})$
$\leq\frac{1}{4}.$
The conclusion then follows from Lemma 4.3. ∎
Finally, we are now at the stage of proving a probabilistic upper bound on the third summand in the error decomposition ($\clubsuit$3.2). The following theorem formalizes this bound.
Theorem 4.14.
Let odd prime $N$ and odd natural number $R$ be sufficiently large such that
$\varepsilon_{2}:=\frac{|\mathcal{A}_{d}(N/2)|}{2}\left(\frac{1}{2}\right)^{R/2}+\frac{N}{2}\left(\frac{3}{4}\right)^{R/2}<1.$ (22)
Then, for any $f\in\mathcal{H}_{d,\alpha,\boldsymbol{\gamma}}$, with probability at least $1-\varepsilon_{2}$,
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq\frac{16\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N_{*}^{2\alpha}}\left(\frac{62N}{15(N-1)}+\frac{2N-1}{N-1}\max\left(960,\sqrt{240N^{4\alpha}_{*}\bar{r}_{\mathrm{sum}}}\right)\right).$
Proof.
Applying Corollary 4.12 to all $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)\setminus K^{\prime}_{N}$ and Lemma 4.13 to all $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}$ satisfying (20), a union bound argument shows that, with probability at least $1-\varepsilon_{2}$, we have:
  • For all $\boldsymbol{h}\notin K^{\prime}_{N}$,
    $\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\delta^{\prime\prime 2}.$
  • For all $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}$ satisfying (20),
    $\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}^{\prime})|^{2}>\frac{F(\boldsymbol{h}^{\prime})}{2}.$
Now, for any $\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}$, either
$|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}<\bar{S}(\boldsymbol{h})^{2}\max\left(960N^{-2\alpha}_{*},\sqrt{240\bar{r}_{\mathrm{sum}}}\right)$
or, if this is not the case, we argue as follows: for any $\boldsymbol{h}\in K_{N}\setminus K^{\prime}_{N}$, by the definition of $K_{N}$,
$\frac{F(\boldsymbol{h}^{\prime})}{2}<\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h}^{\prime})|^{2}\leq\operatorname*{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}\leq\delta^{\prime\prime 2},$
which, together with the definition of $F(\boldsymbol{h}^{\prime})$, implies
$|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq F(\boldsymbol{h}^{\prime})\leq 2\delta^{\prime\prime 2}.$
Combining the upper bounds in the two cases, we obtain
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}|\widehat{f}(\boldsymbol{h}^{\prime})|^{2}\leq|K^{\prime}_{N}\setminus K_{N}|\cdot 2\delta^{\prime\prime 2}+\max\left(960N^{-2\alpha}_{*},\sqrt{240\bar{r}_{\mathrm{sum}}}\right)\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}\bar{S}^{2}(\boldsymbol{h}^{\prime}).$
Finally, using $|K^{\prime}_{N}\setminus K_{N}|\leq N$ and
$\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}\bar{S}^{2}(\boldsymbol{h}^{\prime})$ $\leq 16\left(|K^{\prime}_{N}\setminus K_{N}|\frac{\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\sum_{\boldsymbol{h}^{\prime}\in K^{\prime}_{N}\setminus K_{N}}\sum_{\boldsymbol{\ell}\in N\cdot\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})|^{2}r_{2\alpha,\boldsymbol{\gamma}}(\boldsymbol{h}^{\prime}+\boldsymbol{\ell})\right)$
$\leq 16\left(\frac{N\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}}{N-1}+\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}\right),$
the claim follows. ∎
5 Discussion
In this work, we have extended the median lattice-based algorithm from [22] to the setting where no prior knowledge of the smoothness $\alpha$ and the weight parameters $\boldsymbol{\gamma}$ of the Korobov space is available. A major limitation of our method, as pointed out in Remark 3.3, lies in Step 2 of Algorithm 3.1, where we need to order $\mathrm{median}_{r\in\{1{:}R\}}|\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})|^{2}$ for $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$ and select the $N$ largest frequencies $K_{N}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\}$. A naive implementation requires computing $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$ for all $r\in\{1{:}R\}$ and for every $\boldsymbol{h}\in\mathcal{A}_{d}(N/2)$, whose cardinality grows exponentially with $d$. It is therefore natural to ask whether $K_{N}$, or an adequate substitute, can be found in computation time polynomial in $d$. For example, a mild assumption on $\alpha$ and $\boldsymbol{\gamma}$ may suffice to identify a polynomial-size subset $\mathcal{A}^{*}\subseteq\mathcal{A}_{d}(N/2)$ that contains $K_{N}$. In the extreme case where $\alpha$ and $\boldsymbol{\gamma}$ are fully known, $K_{N}$ can be replaced by the set $\mathcal{A}_{d}(N_{*})$ defined in [22] while still achieving a near $O(M^{-\alpha})$ convergence rate. Another possible approach is to estimate $\alpha$ and $\boldsymbol{\gamma}$ empirically by computing $\widehat{f}_{N,\boldsymbol{z}_{r},\boldsymbol{\Delta}_{r}}(\boldsymbol{h})$ over a small test set for $r\in\{1{:}R\}$. We leave this question for future research.
Acknowledgments
The first and third authors acknowledge the support of the Austrian Science Fund (FWF) Project P 34808/Grant DOI: 10.55776/P34808. The second author acknowledges the support of JSPS KAKENHI Grant Number 23K03210. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. Moreover, the authors would like to thank D. Krieg for valuable comments.